find the slope-intercept form of a line with slope -3 containing point (8,9).
Answer by mananth(16949)

m=-3

Plug value of the slope and point (8, 9) in
Y =mx+b
9.00=-24+b
b=9-(-24)
b=33
So the equation will be
Y =-3x+ 33